OpenStudy (keana):

If the sin of angle x is 4/5, and the triangle was dilated to be two times as big as the original, what would be the value of the sin of x for the dilated triangle?

OpenStudy (perl):

so we multiply each side by 2

OpenStudy (perl):

sin x = 8/10 = 4/5 so it doesn't change the sine value

OpenStudy (keana):

oh ok

OpenStudy (perl):